Determine whether the following series converge, converge absolutely, converge conditionally, or diverge.

  1. "\\displaystyle\\sum_{k=1}^\t\u221e" 3k/k2*2k
  2. "\\displaystyle\\sum_{k=1}^\t\u221e" (-1)k+1 / ( √(k+1) -√k)
  3. "\\displaystyle\\sum_{k=1}^\t\u221e" (-1)k+1 [√(k+1) -√k)]
1
Expert's answer
2021-11-22T17:40:25-0500

Let us determine whether the following series converge, converge absolutely, converge conditionally, or diverge.


1) "\displaystyle\sum_{k=1}^{\infty}\frac{3^k}{k^22^k}" 


Let us use D'Alembert's Ratio Test.


Since "\lim\limits_{n\to\infty}(\frac{3^{k+1}}{(k+1)^22^{k+1}}:\frac{3^k}{k^22^k})
=\lim\limits_{n\to\infty}\frac{3k^2}{2(k+1)^2}=\lim\limits_{n\to\infty}\frac{3k^2}{2(k^2+2k+1)}=\frac{3}2>1,"


we conclude that this series diverge.



2) Since

"\displaystyle\sum_{k=1}^{\infty}\frac{(-1)^{k+1}}{ \sqrt{k+1} -\sqrt{k}}
=\displaystyle\sum_{k=1}^{\infty}\frac{(-1)^{k+1}(\sqrt{k+1} +\sqrt{k})}{ (\sqrt{k+1} -\sqrt{k})(\sqrt{k+1} -\sqrt{k})}
=\displaystyle\sum_{k=1}^{\infty}\frac{(-1)^{k+1}(\sqrt{k+1} +\sqrt{k})}{ ((k+1)-k)}
=\displaystyle\sum_{k=1}^{\infty}(-1)^{k+1}(\sqrt{k+1} +\sqrt{k})," 

and "\lim\limits_{k\to\infty}(\sqrt{k+1} +\sqrt{k})=\infty,"

we conclude that the necessary condition of covergence is not met, and hence this series is divergent.



3) "\displaystyle\sum_{k=1}^{\infty}(-1)^{k+1}( \sqrt{k+1} -\sqrt{k})"


Consider the series "\displaystyle\sum_{k=1}^{\infty}|(-1)^{k+1}( \sqrt{k+1} -\sqrt{k})|=\displaystyle\sum_{k=1}^{\infty}(\sqrt{k+1} -\sqrt{k})."

Since "A_n=(\sqrt{2} -1)+(\sqrt{3} -\sqrt{2})+(\sqrt{4} -\sqrt{3})+\ldots +(\sqrt{n+1} -\sqrt{n})=\sqrt{n+1} -1,"

and "\lim\limits_{n\to\infty}A_n=\lim\limits_{n\to\infty}(\sqrt{n+1} -1)=\infty," we conclude that the series "\displaystyle\sum_{k=1}^{\infty}(-1)^{k+1}( \sqrt{k+1} -\sqrt{k})" diverge absolutely.